OpenStudy (anonymous):

You are going to design an advertisement for a new polynomial identity that you are going to invent. Your goal for this activity is to demonstrate the proof of your polynomial identity through an algebraic proof and a numerical proof in an engaging way! Make it so the whole world wants to purchase your polynomial identity and can't imagine living without it! You may do this by making a flier, a newspaper or magazine advertisement, making an infomercial video or audio recording, or designing a visual presentation for investors through a flowchart or PowerPoint.

OpenStudy (anonymous):

You must: Label and display your new polynomial identity Prove that it is true through an algebraic proof, identifying each step Demonstrate that your polynomial identity works on numerical relationships WARNING! No identities used in the lesson may be submitted. Create your own. See what happens when different binomials or trinomials are combined. Below is a list of some sample factors you may use to help develop your own identity. (x – y) (x + y) (y + x) (y – x) (x + a) (y + b) (x2 + 2xy + y2) (x2 – 2xy + y2) (ax + b) (cy + d)

OpenStudy (anonymous):

these are the identities i have learned in the lesson was the difference of squares, the square of binomials, the difference of cubes, and the sum of cubes

OpenStudy (anonymous):

all im really asking is to help come up with an identity

OpenStudy (anonymous):

@HelpBlahBlahBlah @arabpride @csnider1998 @heather040200 @undeadknight26 @woodaa123 @lilkittykat36 @francoisderozay @skittlesyum @LilVeroDez13 @bean456843 @Vallarylynn16 @damien

OpenStudy (anonymous):

somebody please help

OpenStudy (anonymous):

sorry im horrible at algebra

OpenStudy (anonymous):

its ok at least your being honest

OpenStudy (helpblahblahblah):

sorry i havent learned algebra yet

OpenStudy (anonymous):

oh ok well thats ok

OpenStudy (anonymous):

somebody that knows how to do algebra can yall help

OpenStudy (anonymous):

@Kstate @shamil98 @FATBOI @AngelCriner @xoxoebnap_ @Barrelracer011 @heather040200 @hba @Kittta @Objection

OpenStudy (anonymous):

They have given a bunch of samples. Each one is a polynomial. They want you to combine one or more polynomials and come up with an equation. Here I have chosen to create a polynomial by multiplying (x+a) and (x+a) (Note: (x+a) is one of the samples listed.) (x+a)(x+a) = x^2 + ax + ax + a^2 = x^2 + 2ax + a^2 The left hand side is (x+a)^2. Therefore the polynomial identity is: (x+a)^2 = x^2 + 2ax + b^2 They are asking for a whole lot of other things, but this is the gist of the math part.

OpenStudy (anonymous):

ik what to do i just need help to create a new identity instead of using the ones that i have listed

OpenStudy (anonymous):

i dk

OpenStudy (anonymous):

@xoxoebnap_

OpenStudy (anonymous):

What about x+a^2 ?

OpenStudy (anonymous):

you know what i think the example that you have gave me should be good i really appreciate you helping me @xoxoebnap_

OpenStudy (anonymous):

mhm. medal please?

OpenStudy (anonymous):

and that b should be an a..my typo.

OpenStudy (anonymous):

oh ok thanks